Chuck throws a rock with an initial upward velocity of 12m/sec. How high is it after 0.7 seconds?
Found 2 solutions by nerdybill, Greenfinch:
Answer by nerdybill(7384)

Chuck throws a rock with an initial upward velocity of 12m/sec. How high is it after 0.7 seconds?
.
y = (1/2)at^2 + vt
where
'y' is the height
'a' is gravity -9.8 m/sec^2
v is initial velocity
t is time
.
y = (1/2)(-9.8)(0.7)^2 + (12)(0.7)
y = (-4.9)(.49) + 8.4
y = 6 m

Answer by Greenfinch(383)

u = 12
a = -9.8
t = 0.7
s = ?
s = ut + 1/2(at^2)
s = 12x0.7 + 1/2x(-9.8)x0.49
s = 8.4 - 1.2
s = 7.2 m